OpenStudy (anonymous):

In the xy plane, points with coordinates (0,-5) and (6,-2) lie on line l. Line p contains the point with coordinates (-5,0) and is perpendicular to line l What is the x coordinate of the point where lines l and p intersect?

OpenStudy (anonymous):

Answers they gave me were -6,-5,-4,-3,-2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

you need the equation for both lines. then set them equal and you will find where they intersect

OpenStudy (anonymous):

Can you show me how you got to that answer deeprony

OpenStudy (anonymous):

y+5=(-2+5)/6 (x-0) =y+5=x/2-------this is eq.1 the other eq us perpendicular so, m1*m2=-1 m1 is 1/2 so m2 is -2 forming the eq.of line p using the point-slope formula, y-y1=m(x-x1) whre x1,y1 is -5,0 you gat the eq as, y=-2x-10 now setting the two y's equal...find out x. which, for me, came out to be -2

OpenStudy (anonymous):

*get

OpenStudy (anonymous):

Thanks :]

OpenStudy (anonymous):

ur welkm